I need help with these particular questions on optics. Also if you would please work them out

A equilateral prism (all three angles are 60.0 deg
Answer #1

4. An object (arrow) is placed at a position p away from a thin converging lens. Use a ruler and a ray-trace diagram on this

answers

(a) see rays diagram

(b) from the rays diagram M = -1.5

(c) real

(d) inverted

(e)

\frac{1}{p}+\frac{1}{q}=\frac{1}{f} \Rightarrow \frac{1}{22cm}+\frac{1}{q}=\frac{1}{14cm}

q=38.5cm

so:

M = -\frac{q}{p} = -\frac{38.5cm}{22cm}

\boxed{M=-1.75}  exceeding the value from rays diagram in 16%.